OpenStudy (anonymous):

3 1/4- 6/6= my daughter has this for her homework and I cant remember how to do it! HELP!

OpenStudy (anonymous):

it is easy just do 3*4 then add 1 then find a common denominator

OpenStudy (anonymous):

ok so first you have to make 3(1/4) into a as I like to call them, "top heavy fraction" multiply 4*3=12+1=13/4 13/4-6/6

OpenStudy (anonymous):

FInd a common denominator which is twelve because both 6 and 4 go into 12

OpenStudy (anonymous):

13/4=39/12 because: 4*3=12 and 13*3=39

OpenStudy (anonymous):

3 1/4 -6/6 =___ 6/6 is six devided by siz which = 1 so 3 1/4 -1= 2 1/4 or 9/4

OpenStudy (anonymous):

Now do the same thing for 6/6 6/6=12/12 because: 6*2=12 6*2=12

OpenStudy (anonymous):

i said it is easy

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

your problem should now look like this: 39/12-12/12=? you can simply subtract across the top now to find your answer your answer should be: 27/12

OpenStudy (anonymous):

why the heck are people using common denominators? 6/6 =1 why not reduce first?

OpenStudy (anonymous):

that answer is 2 1/3

OpenStudy (anonymous):

2 1/4

OpenStudy (anonymous):

that is what i meant

OpenStudy (anonymous):

you can simplify this by asking how many times does 12 go into 27 Answer: 2 times and how much is left over 2(3/12) Your final answer should be : 2(1/4)

OpenStudy (radar):

Bethany105 I would look at a problem and try to make it simple. In this case I would look at the fraction this being subtracted 6/6 hey! thats a 1! Now your problem is 3 1/4 - 1 or just plain ole 2 1/4